OpenStudy (anonymous):

Approximate the real number solution(s) to the polynomial function f(x) = x3 + 2x^2 – 5x – 6 .

OpenStudy (anonymous):

\[\huge f(x) = x3 + 2x2 – 5x – 6 = (x-2)(x+1)(x+3) \] roots are -3 -1 2